Find the minimum and maximum values of Y in 4X^2+12XY+10Y^2-4Y+3??Ans:1,3 res

Grade:

1 Answers

Arun
6 years ago
Dear Aayushi
 
=1+ (y-2)2(2x + 3y)- 1+ (y-2)2= (2x + 3y)– 4y + 32+ 12xy + 10 y2Let f(x) = 4xThanks for your post. The question is a bit unclear, I am assuming you want to know the minimum and maximum values of the whole function (as it makes sense)Dear student, |y- 2|  1 Hencemaximum value of f(x) = 3Therefore , minimum value of f(x) = 1
 
 
 
Regards
Arun (askIITians forum expert)